A construction firm manufactures steel and is normally distributed with mean equal to 800MPa and a standard deviation of 40 MPa before it reaches yielding. Find the probability that the steel is in between 778 MPa and 834 MPa.


1
Expert's answer
2021-05-31T06:09:45-0400

"P(778<X<834)=P[\frac{778-800}{40}<\frac{X-800}{40}<\frac{834-800}{40}]"

"=P(-0.55<Z<0.85)"

"=P(Z<0.85)-P(Z<-0.55)"

"=0.80234-0.29116"

"=0.51118"